OpenStudy (anonymous):

a larger cube has a volume of 64. The side of a smaller cube is 1/4 that of the larger cube what is the volume of the smaller cube?

OpenStudy (anonymous):

Gosh i need help someone please answer.

OpenStudy (anonymous):

Volume:\[V=length*width*height\] For a cube, all the side lengths are the same, so plug in "s" for the length, width, and height: \[V_{large}=s_{large}*s_{large}*s_{large}=s_{large}^3=64\] Solve for the large s: \[s _{large}=\sqrt[3]{64}=4\] If the side of the smaller cube is 1/4 that of the larger, you have: \[s_{large}=4s_{small}\] We know the large s is 4, so plugging it in:\[4=4s _{small}\] Therefore: \[s _{small}=1\] Plug the small s into the volume equation:\[V _{small}=s _{small}^{3}=(1)^3=1\]

OpenStudy (anonymous):

Thank you soooo much Chris.

OpenStudy (anonymous):

Wait... i'm looking for volume not surface area.